OpenStudy (anonymous):

How do you solve this sorta equations: b^x = ax? Specifically, this example : (1.09)^x = 0.546x

OpenStudy (anonymous):

damn it's different !

OpenStudy (anonymous):

i think we have to equal the base then when base is equal we can just ignore the bases so we will try to solve the exponents for x for exmple: 25^x=5^x-1 we can write 25 like this 5^2 so we will write instate of 25: 5^2x=5^x-1 now because the bases are equal so we will ignore them we will solve the exponent for x 2x=x-1 x=-1 hope i could help you

OpenStudy (anonymous):

@Roh its b^x = ax and not b^x = a^x

OpenStudy (anonymous):

@mridul1 so dear b^x=ax so we will get that x=1 then we will solve it ..

OpenStudy (anonymous):

Look in Examples section here: http://en.wikipedia.org/wiki/Lambert_W_function

OpenStudy (anonymous):

hey thanks people. looks complicated indeed. is there any easy way to evaluate the W function? In this case, x = -W(-1.67) / ln(2) .. now?

OpenStudy (anonymous):

Use the Taylor series expansion on the same page (or Wolfram).

OpenStudy (anonymous):

right! thanks dude.. you rock

OpenStudy (anonymous):

ur welcome.